OpenStudy (anonymous):

Simplify the expression 28x -17 +12x + 13. Show your work.

OpenStudy (lilg132):

rearrange 28x + 12x - 17 + 13 = 40x - 4

OpenStudy (anonymous):

bring like terms together 28x + 12x - 17 + 13 = 40x - 4

OpenStudy (anonymous):

can you answer ?Explain why the expression 7x + 5xy2 – 2x is a binomial when it is simplified

OpenStudy (anonymous):

7x - 2x = 5x so this simplifies to 2 terms: 5xy^2 + 5x which is a binomial ( consists of two terms)